OpenStudy (anonymous):

How do i graph this; x-y = -2

OpenStudy (shamil98):

-y= -x -2 y=x+2 2 is the coordinate the slope is 1/1

OpenStudy (anonymous):

when manipulating an equation, do the same thing to both sides. For this equation, add y to both sides, then add 2 to both sides.

OpenStudy (anonymous):

that looks like two points

OpenStudy (anonymous):

oh .... awkward

OpenStudy (anonymous):

Move y's to one side, and move the rest to other side. x-y=-2 x-y+y=-2+y x=-2+y x+2=-2+y+2 x+2=y

OpenStudy (anonymous):

Now, you can see that it is a linear function of slope m=1, y intercept 2.